OpenStudy (anonymous):

Find real numbers a, b, and c so that the graph of the quadratic function y=ax^2+bx+c contains the points (-2, 13), (0,1), and (-2,3)

OpenStudy (anonymous):

Might have to solve a system of linear equations here. You cool with that?

OpenStudy (anonymous):

yes! i've solved some of this already, but it seems inconsistent to me once I build the second set of equations. I'm ready! :)

OpenStudy (anonymous):

Ok. Set up the equations, you have three of them, but one will dissolve quite nicely since it is the y-intercept. y=ax^2+bx+c (-2, 13): 13=a(-2)^2+b(-2)+c (0,1): 1=a(0)^2+b(0)+c (-2,3): 3=a(-2)^2+b(-2)+c Hmm, yeah, I see what you're saying about being inconsistent. Are you sure you have the points right? There shouldn't be two y-values for a single x input..

OpenStudy (anonymous):

Maybe this is supposed to be x=ay^2+by+c? Have you done general quadratics and conic sections yet? This could be a 'sideways' parabola.

OpenStudy (anonymous):

i attached the actual page for you to view

OpenStudy (anonymous):

I can't open Powerpoint files (especially ones with embedded macros - my antivirus software will throw a fit over that).

OpenStudy (anonymous):

ok, in that case, yes, i'm sure that's the exact problem. word for word.

OpenStudy (anonymous):

i got it down to the original 3 equations. 4a-2b+c=13....c=1....4a-2b+c=3. That's where i multiplied them through, eliminating c, and ended up with 4a-2b=12 and 4a-2b=2. Is my process up to that point correct? Because that's when I end up with 0=10 and just can't continue. Is there an approach on the calculator I could utilize? with the matrix option?

OpenStudy (anonymous):

I don't think matrices are necessary. It might be a degenerate quadratic that reduces to a single straight line, but the points don't seem to add up. I'll check it again.

OpenStudy (anonymous):

parallel perhaps? that would support the inconsistent status atleast

OpenStudy (anonymous):

so it's not just me.....phew!

OpenStudy (anonymous):

Yeah, the system reduces to 2a-b=6 2a-b=1 which is inconsistent for a and b, but we can at least trust c=1. Maybe the equation is simply y=1 and is a horizontal line.

OpenStudy (anonymous):

It doesn't work for the other two points though, which means it can't be a function of x. It could be a function of y, though. Try the sideways parabola.

OpenStudy (anonymous):

i'm not sure what you mean. we just learned this on wednesday, new material, but this was the problem our teacher challenged us to over the weekend. hmmm

OpenStudy (anonymous):

we begin matrices on monday, does that guide you at all here?

OpenStudy (anonymous):

Nope. You can use matrices if you want, but it will yield exactly the same results. The system reduces to a 2x2, which is easy to do with matrices (especially with the Cramer's Rule short cut, but is completely unnecessary). Try graphing the thing. Plot the points and try drawing a parabola through them.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

eh, it's more like half of a x^2+1graph, what does that mean? definitely not going to be sideways.

OpenStudy (anonymous):

Ah, then you can ignore one of the points and only use the first quadrant part.

OpenStudy (anonymous):

wouldn't it be quadrant II?

OpenStudy (anonymous):

Yeah, had to use my imagination a little more. Here's the sideways parabola idea: